OpenStudy (caylie00):

Okay this one is a little hard for me but I might be wrong. |a|-3/4=-5/8 A.1/8 or -1/8 B.7/8 or -7/8 C.1 3/8 or -1 3/8 D. none

OpenStudy (anonymous):

move -3/4 to the other side of the = sign to get the absolute value of a alone

OpenStudy (anonymous):

first, we add 3/4 to both side \(|a|-\dfrac{3}{4}=-\dfrac{5}{8}\\ |a|=-\dfrac{5}{8}+\dfrac{3}{4} \\|a|=-\dfrac{5}{8}+\dfrac{6}{8}\)

OpenStudy (caylie00):

Okay thank you I got the hang of it now. I can finish it :)

OpenStudy (anonymous):

post your answer if you want it confirmed

OpenStudy (caylie00):

Okay my answer is A.1/8 or -1/8

OpenStudy (anonymous):

well done :)

OpenStudy (caylie00):

Why thank you!